OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

We know that parallel lines have the same slope. What is the slope of your given line? (Easier to see if you solve for y first)

OpenStudy (anonymous):

or, \[-\frac{ 4 }{3 }\]

OpenStudy (anonymous):

Solve for y first. \[4x - 3y = -12\] \[-3y = -4x - 12\] \[y = \frac{4}{3}x + 4\]

OpenStudy (anonymous):

So, we need a line that has the same slope (4/3), and passes through the given point. We can just put this information into the point-slope form for a line: \[y-y_o = m(x-x_o)\]

OpenStudy (anonymous):

My choises are: \[Y = \frac{ 4 }{ 3 }x + 8\] \[Y = \frac{ 4 }{ 3 }x + 3\] \[Y = -\frac{ 4 }{ 3 }x + 8\] \[Y = -\frac{ 4 }{ 3 }x + 3\]

OpenStudy (anonymous):

Just plug in what you know and solve for y. It'll work out :)

OpenStudy (anonymous):

I think its B?

OpenStudy (anonymous):

Functions just lose me. :'(

OpenStudy (anonymous):

\[y - 4 = \frac{4}{3}(x+3)\] \[y-4 = \frac{4}{3}x + 4\] \[y = \frac{4}{3}x + 8\]

OpenStudy (anonymous):

So its A?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you.